### v1

# A Comprehensive Survey of Foundation Models in Medicine

Abstract
Foundation models (FMs) are large-scale deep-learning models trained on extensive datasets using self-supervised techniques. These models serve as a base for various downstream tasks, including healthcare. FMs have been adopted with great success across various domains within healthcare, including natural language processing (NLP), computer vision, graph learning, biology, and omics. Existing healthcare-based surveys have not yet included all of these domains. Therefore, this survey provides a comprehensive overview of FMs in healthcare. We focus on the history, learning strategies, flagship models, applications, and challenges of FMs. We explore how FMs such as the BERT and GPT families are reshaping various healthcare domains, including clinical large language models, medical image analysis, and omics data. Furthermore, we provide a detailed taxonomy of healthcare applications facilitated by FMs, such as clinical NLP, medical computer vision, graph learning, and other biology-related tasks. Despite the promising opportunities FMs provide, they also have several associated challenges, which are explained in detail. We also outline potential future directions to provide researchers and practitioners with insights into the potential and limitations of FMs in healthcare to advance their deployment and mitigate associated risks.
1 Introduction
1.1 History
Artificial intelligence (AI) has achieved remarkable milestones in recent decades. This progress is mainly attributed to the enhanced availability of large-scale data, computational resources, and intelligent algorithm design [1]. AI-based techniques have been used for decades within healthcare [2]. For example, computer-assisted surgery began in the early 1980s, which marked the start of medical AI [3]. In 1989, a small convolutional neural network (CNN) model demonstrated promising performance by using a backpropagation algorithm to recognize handwritten digits [4]. However, the application of AI in medicine was not facilitated until the AI spring, which was initiated by the emergence of the AlexNet model [5]. The significant momentum in the development of AI models was demonstrated after the arrival of the AlexNet model. This was trained on the most prominent training dataset called ImageNet [6]. Subsequently, other state-of-the-art deep learning (DL) models have been developed to surpass AlexNet [5] through the ImageNet Large Scale Visual Recognition Challenge (ILSVRC) [6]. Following the CNN, Recurrent neural networks (RNN) [7] were proposed for sequential data processing that showed promising performance in various domains including time series analysis, natural language processing (NLP), and speech recognition [8]. Another significant contribution came with the introduction of generative adversarial networks (GAN) [9] for generating synthetic data [9, 10].
Transformers, introduced in 2017, formed a crucial foundation of the most powerful deep learning models available today. Transformers introduced a self-attention mechanism to capture sequences in parallel and addressed the limitations of computational and sequential RNNs. Various state-of-the-art models have been developed based on the success of transformers. Transformer models have also played a key role in driving the creation of foundation models (FM) [11]. FMs are large-scale pre-trained neural networks trained on large amounts of data that can be fine-tuned to perform a wide range of downstream tasks. The key enablers of FMs include training data, base model, transfer learning, and scale [11]. FMs adopt a self-supervised learning (SSL) approach, unlike traditional DL models that require a large amount of data. In this approach, the pre-training task is autonomously generated from unlabeled data. SSL evolved from word embeddings that provided context-independent vectors for words [12, 13] to autoregressive language modeling that predicts the next word based on previous words [14, 15].
SSL led to one of the first FMs known as BERT [16]. A powerful transformer architecture characterized by its bidirectional sentence encoders and scalability extending to larger architectures and datasets. BERT became the standard model in 2019. Many subsequent models were developed based on BERT architectures, including RoBERTa [17], BART [18], and T5 [19]. These models have become prevalent in the development of state-of-the-art NLP models. These NLP models are termed large language models (LLM). Despite being available since late 2018, these models gained broader recognition across various domains following the introduction of ChatGPT in late 2022111https://chat.openai.com/. Furthermore, they were soon generalized to different fields, such as text [20], images [21], videos [21], speech [22], tabular data [23], protein sequences [24], organic molecules [25], and reinforcement learning [26].
Healthcare constitutes a substantial economic sector representing 17% of the United States GDP [11]. Many specialized models have been developed for various healthcare tasks [27]. However, they are based on a limited amount of labeled data and face challenges when generalized to different tasks. It remains challenging to adopt general-purpose FMs in the healthcare domain due to the complex structures and limited publicly available medical data. Therefore, several works have been carried out to define FMs specifically for medical data. These models are task-agnostic and can be generalized across different tasks within the medical domain. For instance, the GatorTronGPT [28] model was built on the GPT-3 [29] architecture and includes 20 billion parameters. It was trained on 277 billion words with 82 billion words of clinical text which were obtained from 2 million patients. GatorTronGPT exhibits strong performance in various medical tasks such as named entity recognition (NER), biomedical question answering (QA), biomedical relation extraction, natural language inference (NLI), and semantic textual similarity. Similarly, models such as MedSAM [30] trained on 1.5 million medical images, are used for medical image segmentation and surpass other models [31, 32]. FMs designed for other tasks within the medical domain are described in detail in Section 5.
1.2 Research Objectives and Outline
This survey paper aims to provide a comprehensive review of FMs in the healthcare domain. We first provide an overview of the historical progression of artificial intelligence (AI) and deep learning (DL) models followed by the development of transformers-based models (shown in Figure 1). The survey is structured to provide a comprehensive study of FMs, with Section 2 explaining the background of FMs, followed by an extensive overview of FMs in Section 3. The flagship FMs such as the BERT family, GPT models, CLIP, stable diffusion, and industrially scaled pre-trained models are explained in Section 4. The core FMs specifically designed for medical applications are presented in Section 5. The applications of FMs in the healthcare domain are presented in Section 6. We present multiple applications including clinical NLP (Section 6.1), medical computer vision (Section 6.2), healthcare graph learning (Section 6.3), biology and omics (Section 6.4), and others (Section 6.5). In Section 7, we explain the available opportunities associated with FMs, including controllability, adaptability, and applicability. The challenges associated with FMs, such as cost, interpretability, validation, and scale, are thoroughly discussed in Section 8. A discussion and potential future work are presented in Section 9 and a conclusion is provided in Section 10.
Figure 1: Outline of the proposed work.
1.3 Scope
FMs have been widely adopted across multiple domains. However, a detailed taxonomy covering all aspects of FMs within the medical domain is still missing. Existing surveys have concentrated on general-purpose FMs [11, 33, 34, 35, 36], whereas this survey focuses on the healthcare domain. Furthermore, we describe the major flagship models that led to the development of medical-based FMs (Section 4).
1.4 Existing surveys
In 2023, the rapid advances in FMs led to the publication of multiple survey articles (Table 1). However, the majority of these articles do not focus on healthcare. Some of the articles only focus on LLMs on [37], which excluded FMs for other modalities such as diffusion models [38]. Furthermore, existing articles focused on healthcare lack a detailed taxonomy of FMs that includes all of its associated aspects, such as LLMs, vision, protein, audio, and graph FMs.
Table 1: Comparison with existing survey.
Work Healthcare LLMs Vision FMs Protein Audio Graph FMs
Wornow et al. [39]
Zhou et al. [33]
Yang et al. [38]
Yang et al. [34]
Awais et al. [35]
Bommasani et al. [11]
Kolides et al. [36]
Hadi et al. [40]
Azad et al. [41]
Zhou et al. [42]
Zhao et al. [43]
This survey
1.5 Contribution
This survey provides a comprehensive analysis of recent works published until early 2024. It encompasses the evolutionary journey of FMs in the healthcare sector, providing insights into their historical development and categorization. The survey explores the applications of FMs in various healthcare domains including clinical NLP, medical computer vision, biology and omics, and other relevant areas. Furthermore, we highlight the major challenges associated with FMs and outline future directions associated with FMs in healthcare.
2 Background
2.1 Transformer
A Transformer [44] is a neural network architecture initially developed for sequence modeling. It addressed the inefficient computation and poor performance of RNNs caused by the sequential processing of the information. Due to the attention mechanism in transformers, sequences can be processed in parallel, which is faster and more efficient during training. The attention mechanism in a transformer is computed by taking a weighted sum of the representations of all tokens in a sequence, where the weights are determined by the similarity between the tokens. Since there is no convolution or recurrence, positional encoding is used to represent the relative positions of tokens in a sequence. The positional encoding vectors are added to the embedding vectors for the tokens at the input to the self-attention mechanism. This helps with learning long-range dependencies in a sequence by attending to the positional encoding vectors of the tokens. Transformers have led to the development of state-of-the-art models including BERT [16], GPT [15][45], and other models [35]. Transformers were originally developed for NLP tasks, but they have since been extended to other domains, including computer vision. Vision transformer (ViT), first proposed in 2021 [46] are a type of transformer specifically designed for computer vision tasks. It converts an image into a sequence of patches and then calculates meaningful embedding using a self-attention mechanism. The positions of patch embeddings are retained using positional encoding.
2.2 Attention
Attention mechanism in deep learning originated from the human visual attention that selectively focuses on one or more aspects of information within a vast array of available data [47]. For instance, Itti et al. [48] proposed visual attention, inspired by the primate visual system, to identify salient regions. The attention mechanism has been used mainly with RNNs or CNNs [49] for multiple tasks [47]. However, transformers rely solely on attention, unlike CNNs and RNNs which use attention in part of the model. The attention mechanism in transformers computes the similarity between tokens in a sequence. The attention function maps a query and a set of key-value pairs to an output, where the output is a weighted sum of the values, with the weights determined by the similarity between the query and each key [44]. Multiple attention mechanisms have been proposed recently, including self-attention [44], multi-head attention [44], masked attention [50], and criss-cross attention mechanisms [51]. In self-attention, each element in the input sequence is associated with query, key, and value vectors. The model then computes attention weights by measuring the similarity between the query and key vectors for all elements in the sequence. These weights guide the model in taking a weighted sum of the value vectors for each element, effectively enabling the model to selectively focus on different parts of the input when processing each element. Multi-head attention learns multiple representations of the input sequence, each of which can attend to different aspects of the sequence. This is done by performing multiple self-attention operations in parallel, on different input representations. Cheng et al. [50] proposed Mask2Former which utilizes masked attention in the transformer decoder. Masked attention restricts attention to specific regions of interest, such as the foreground region of a predicted mask, during the calculation of attention scores. Criss-cross attention [51] efficiently captures full-image contextual information for tasks like semantic segmentation and object detection. It utilizes a novel Criss-Cross Network with a criss-cross attention module, allowing each pixel to gather contextual information from its surroundings.
2.3 Self-supervised learning (SSL)
The limitation of supervised learning is that it depends heavily on large amounts of labeled data, which is a time-consuming, expensive, and often infeasible task [52] [53]. Furthermore, since the data are annotated based on the labels due to correlation, it introduces generalization error [52] and is prone to adversarial attacks [53]. In contrast, self-supervised learning (SSL) can be used to train the model on large amounts of data without labels. SLL leverages the inherent structure or context within the data itself to create meaningful representations. SSL can be divided into three main categories: contrastive, generative, and adversarial learning [53]. Contrastive learning algorithms train models to distinguish between similar and dissimilar data points based on similarity within the representation space. Generative learning algorithms train models to generate realistic data. Adversarial learning is a hybrid generative-contrastive learning approach that improves representations by reconstructing the original data distribution rather than individual data samples, using discriminative loss functions as objectives.
2.4 Human feedback reinforcement learning
Reinforcement learning (RL) is an ML-based algorithm in which an agent interacts with an environment and learns to make a sequence of decisions to maximize a cumulative reward [54]. It can be formulated using a Markov Decision Process (MDP), mathematically, $\langle S,A,T,R,\gamma\rangle$ [54, 55], where $S$ represents multiple states in an environment and selects actions ($A$). The transition probability ($T$) specifies how the agent moves from one state ($s$) to another ($s^{\prime}$), and a reward function ($R$) provides feedback on the actions taken. The discount factor ($\gamma$) is a critical parameter that influences the relative importance of future discounted rewards in the decision-making process. The agent’s goal is to find a policy ($\pi:S\rightarrow A$) that specifies how to select actions in each state to maximize the expected reward. Although RL learns complex behavior with predefined reward functions, however, it can be challenging to express some behaviors as rewards [55]. Therefore, human feedback RL (HFRL) in which the agent learns about their action from feedback from a human trainer. The human feedback has resulted in significant performance improvements [56].
2.5 CLIP
Traditional DL algorithms have demonstrated promising performance in multiple computer vision tasks. However, they rely heavily on extensive labeled training data, which is both labor-intensive and expensive to obtain. Additionally, these algorithms struggle to generalize beyond the specific tasks they are trained for [57, 58, 59]. These challenges are addressed by introducing CLIP [57] which is a novel neural network-based architecture that leverages natural language-supervised vision data readily available on the internet, effectively reducing the dependence on expensive labeled datasets while enabling broader generalization.
CLIP is a robust neural network architecture with zero-shot capabilities [60] that achieves various classification benchmarks using natural language instructions, without direct optimization for each specific benchmark [60]. It is a simplified variant of ConVIRT [61] and is trained on a dataset called WebImageText (WIT), comprising 400 million pairs of images and corresponding text collected from the internet. Multiple works such as [61, 62, 63, 64]utilized text-image representations, however, these methods operate under supervised, weakly-supervised, self-supervised, or unsupervised learning paradigms. In contrast, CLIP utilizes natural language as its training signal. The key advantage of leveraging natural language supervision for learning is its scalability, unlike traditional crowd-sourced labeling which requires classic ML-compatible annotations. CLIP can passively learn from the abundant textual data available on the internet, establishing connections between visual representations and language, which enables flexible zero-shot transfer to a wide range of tasks
CLIP employs separate transformer encoders for text and image inputs. The visual encoder is based on the ViT, while the text encoder utilizes the GPT-2 transformer. To enhance efficiency, CLIP employs a contrastive objective, predicting which text corresponds to a given image. It jointly trains an image and a text encoder to maximize the similarity between embeddings of paired (image, text) examples while minimizing similarity with other combinations. This process results in a multi-modal embedding space where images and their associated texts are closely related. The existing variations of CLIP-based models include CRIS [65], CLIPseg [66], FLIP [67], SLIP [68], MaskCLIP [69], and GLIP [70].
Although CLIP is a robust model trained on large amounts of data, medical image-text datasets are of a much smaller scale than general internet data. Therefore, CLIP can be prone to false negatives, such as incorrectly treating images and reports from separate patients as dissimilar, despite their shared semantic content. Furthermore, it relies on image-text pairs while medical images are provided with medical domain labels compared to raw text. Therefore, Wang et al. [71] proposed MedCLIP, which decouples images and texts for contrastive learning, significantly increasing the usable training data and replacing the InfoNCE loss [72] with a semantic matching loss based on medical knowledge (semantic similarity) to eliminate false negatives. BioClinicalBERT [73]  [^2] https://huggingface.co/emilyalsentzer/Bio_ClinicalBERT has been used as the text encoder and Swin Transformer [74] or ResNet-50 [75] have been used as the vision encoder. Detailed experimental results showed that the proposed MedCLIP achieved promising performance compared to CLIP and other methods [76] [61] with 10 times fewer training data. Zhang et al. [77] proposed BiomedCLIP, which outperformed the original CLIP in the biomedical domain. To address the growing need for large-scale datasets in the field of biomedical research, Lin et al. [78] introduced a substantial dataset comprising more than 1.6 million image-text pairs. The dataset was collected from over 2.4 million scientific papers with over 381K medical figure-caption pairs. Additionally, the authors presented a novel CLIP-based model named PMC-CLIP, pre-trained on the proposed dataset, to support various healthcare applications. Liu et al. [79] proposed a CLIP-based framework that uses text embeddings from CLIP to enhance segmentation models for automated organ segmentation and tumor detection which can efficiently segment 25 organs. However, their proposed method relies significantly on one labeled training dataset. Therefore, Liu et al. [79] proposed a CLIP-based universal model that incorporates text embeddings and masked backpropagation.
3 Overview of Foundation Models
FMs revolutionized the way intelligent algorithms work by enabling a single model to perform various downstream tasks after training on large-scale unlabeled datasets [39]. The fundamental architecture of FMs relies on deep neural networks, particularly transformers trained on large-scale data. Below we provide a brief overview and describe the working mechanism of FMs.
3.1 Learning architecture
The underlying architecture of FMs is typically based on transformers capable of capturing long-range dependencies and contextual relationships within data (Section 2.2). The main architectures are encoder-decoder, encoder only, and decoder only are explained below:
3.1.1 Encoder-decoder
The original transformer model [44] is based on encoder-decoder architecture. The encoder maps the input sequence of symbol representations $X=(x_{1},x_{2},\ldots,x_{n})$ to a embedding space $z=(z_{1},z_{2},\ldots,z_{n})$. Then the decoder works in an auto-regressive manner and generates the output $Y=(y_{1},y_{2},\ldots,y_{n})$ one element at a time. This process involves utilizing the previously generated symbols as additional input at each step, allowing the model to factor in context from the ongoing generation [44]. Multiple FMs use encoder-decoder architecture, including BART [18], Segnet [80], and T5 (Text-to-Text Transfer Transformer) [19].
3.1.2 Encoder-only
Encoder-only models encode contextual information from input sequences that extract and encode relevant rich and informative representations of the input sequence that can be used for various downstream tasks. Encoder-only models include BERT [16], ViT [46], and CLIP [57].
3.1.3 Decoder-only
Decoder-only models specialize in the creative generation of coherent sequences. They predict the next token when given a sequence of tokens, which results in the generation of contextually meaningful outputs. Many state-of-the-art models, such as BLOOM [81], PaLM [82], and open pre-trained transformers [83], are based on decoder-only architecture. GPT [15] models, such as GPT-2 [60], GPT-3 [29], and the more recent GPT-4 [84], are decoder-only models pre-trained on large-scale unsupervised text data.
3.2 Learning algorithms
SSL enables FMs to learn rich contextual representations from a large amount of unlabeled data. The learned representations can then be fine-tuned for specific downstream tasks. SSL tackles the challenges associated with expensive, biased annotation and narrowly applicable supervised learning by providing a more robust and efficient approach for generalizing across downstream tasks [85]. SSL methods that require supervised fine-tuning steps with manually labeled data have limited usability, especially for medical datasets, which are difficult to collect and annotate [86]. Therefore, zero-shot learning-based methods are used instead, which do not require explicit manual or annotated labels during training [86]. Zero-shot learning-based methods include the GPT models [15, 60, 29, 84], CLIP [57], PaLM[82], and SLIP [68].
3.3 Human feedback reinforcement learning
FM models that learn from human feedback tend to demonstrate improved performance and can be generalized to other domains, unlike supervised learning [87]. For instance, Stiennon et al. [87] used the GPT-3 model with 6.7 billion parameters to predict human-preferred summaries from large-scale data and used the model as a reward function to fine-tune a summarization policy through RL. Specifically, for Reddit posts [88], multiple summaries were sampled from various sources including the current policy, initial policy, original reference summaries, and baselines. The summaries were then evaluated by human labelers and a reward model was trained to predict the likelihood that one summary was preferred over the other, based on human judgments.
3.4 Prompting Foundation Models
Prompting FMs are directed to perform a specific task to enhance their effectiveness and relevance [35]. In promoting FMs, users provide a prompt or a series of prompts that guide the model’s response generation process. It aids the model to generate a response that is more aligned with the user’s expectations. Incorporating additional context into the training and evaluation prompting FMs aids in improving the model’s performance by providing relevant information and guiding its decision-making process. For instance, GPT-based application [29, 84] possesses strong language understanding capabilities that can be prompted with specific textual inputs, such as questions. These can be further fine-tuned for tasks like text generation, sentiment analysis, or document summarization. Furthermore, prompts can be used to specify the desired task such as image classification or caption generation that allows the model to focus its attention and resources accordingly. Models such as CLIP [57] which is trained on both image and text encoders using a contrastive pre-training objective rely on textual prompts for supervision. Furthermore, CLIP variants such as MaskCLIP [69] and RegionCLIP [89] are based on object detection and segmentation to align image regions with textual descriptions. We refer readers to [35] for further details about prompting models.
4 Flagship Foundation Models
This section highlights the key foundation models and their variants.
4.1 BERT
BERT [16] is a transformer-based model that was developed for language representation through training on a large corpus of textual data. It is a pre-trained bidirectional transformer model that utilizes contextual information within sequences from unlabeled data through the use of masked language modeling (MLM). The BERT frameworks consist of two main stages: pre-training and fine-tuning. The pretraining stage consists of two tasks: MLM and next sentence prediction (NSP). For MLM, 15% of input tokens are randomly masked, and the model predicts these masked tokens. However, [MASK] tokens are not present during fine-tuning, causing a mismatch between pre-training and fine-tuning phases. To address this, the authors implemented a strategy during pre-training where [MASK] tokens were utilized 80% of the time, and a random token 10% of the time, and the unchanged token was used 10% of the time during pre-training. The authors showed that existing techniques were unable to determine the relationship between multiple sentences and were only focused on individual sentences. Therefore, NSP was used to predict the next sentence that achieved over 97% accuracy for the binary classification task (next sentence vs not next sentence).
During fine-tuning, a self-attention mechanism is used for downstream tasks with a limited amount of labeled data. Originally, there were two BERT variants (BERTBASE and BERTLARGE) distinguished by several transformer blocks, hidden layers, and self-attention heads. BERTBASE had 110M parameters while BERTLARGE consisted of 340M parameters. However, following the success of BERT, multiple variants of BERT models were introduced. For instance, a more efficient and robust model called RoBERTa (robustly optimized BERT pretraining approach) [17] surpassed the BERT model using a few modifications to the original BERT. For example, the RoBERTa is trained for longer (500K steps) on a large dataset of over 160 GB of English-language corpora. Furthermore, the [MASK] token was applied dynamically which creates a new masking pattern each time at the input to the model. In addition, the NSP loss module was removed.
Although BERT has been reasonably successful, its computation complexity and communication overhead are major limitations to adopting such models at scale [90]. Therefore, models inspired by BERT were developed to achieve BERT-level performance while reducing the training time and model size. For instance, Sanh [91] proposed the lightweight DistilBERT, which is a smaller and faster language representation model. DistilBERT is based on the distillation knowledge compression technique [92], which is a smaller model (student) and trained to reproduce a larger model (teacher). The number of layers in DistilBERT is reduced by half and the architecture has been optimized [91]. Furthermore, the triple loss was used during pretraining: masked language modeling loss, cosine embedding loss, and distillation loss. The training objective in distilBERT included the combination of masked language modeling loss, distillation loss, and cosine embedding loss to align student and teacher hidden state vectors. The proposed model was nearly 40% smaller in size and 60% faster than BERT with comparable performance on various downstream tasks.
A Lite BERT (ALBERT) [90] introduced factorized embedding parameterization and cross-layer parameter sharing that incorporated BERT. Furthermore, the NSP module was replaced with a self-supervised loss for sentence order prediction to improve its performance. The proposed model was 1.7 times faster, used 18 times fewer parameters, and achieved better performance than BERT. The use of BERT-based models have been deployed in healthcare for various applications with great success. A model based on a denoising autoencoder known as BART [18] has been developed for sequence-to-sequence modeling and shows promising performance. However, it is computationally expensive, with 10% more parameters than BERT.
4.2 GPT
The ChatGPT we interact with today originated from extensive research trials based on a GPT model. Radford et al. [15] proposed a transformer-based GPT model that utilizes an unsupervised approach during pre-training and supervised fine-tuning for natural language understanding tasks. GPT was trained on the BooksCorpus dataset [93] for unsupervised pre-training. Then, the parameters learned at the pre-training stage were used for a target task with a manually annotated dataset in a supervised manner. A separate work by Radford et al. [60] presented a GPT-2 language model with a zero-shot setting for multiple downstream tasks without the need for any major modifications. GPT-2 was trained on a large amount of data — WebText — that contains over 8 million cleaned documents with more than 40 GB of text. GPT2 performed well on multiple tasks, including reading comprehension, summarization, translation, and question-answering. However, although it performed better than baselines, GPT-2 still could not be applied practically as the performance was still not promising. Therefore, a more robust model, GPT-3, with 175B parameters, was presented and performed well on various downstream tasks. GPT-3 uses the same architectures as GPT-2 except for the implementation of alternating dense and locally banded sparse attention patterns in the transformer layers. It was trained on a large amount of data with 570 GB (300 billion tokens) of plain compressed filtered text. The performance of GPT-3 for various downstream tasks surpasses that of existing techniques when tested using zero-shot, one-shot, and few-shot learning scenarios. One key solution to reduce untruthful or irrelevant outputs is to incorporate human feedback, such as in InstructGPT [94]. InstructGPT uses reinforcement learning from human feedback (RLHF) to fine-tune the GPT-3 model on human data and reduce the number of parameters to only 1.3B. Another more sophisticated language model known as ChatGPT  [^3] https://chat.openai.com/ or GPT3.5 provides more user-friendly, interactive, and context-aware responses and conversational experience. It follows a similar refined architecture to that of InstructGPT [94]. There are a range of healthcare applications that utilize ChatGPT and these will be covered in later sections. ChatGPT is limited to text inputs, but this limitation was addressed in the GPT-4 [84] model, which accepts images in addition to text inputs while producing text outputs. However, the details of GPT4 are not publicly available and a major issue with such language models is hallucinations [84].
4.3 DALL-E
DALL-E [95] is a transformer-based text-to-image generative model that employs a two-stage training procedure. In the first stage, each RGB image is compressed into a 32$\times$32 grid of image tokens using a discrete variational autoencoder (dVAE). It bypasses the direct use of pixels, significantly reducing memory consumption. Furthermore, short-range dependencies between pixels are prioritized during training [96]. Therefore, in the second stage, an autoregressive transformer models the joint distribution over text and image tokens. The training maximizes the evidence lower bound (ELB) on the model’s joint likelihood over images, captions, and tokens. The proposed model was trained on a dataset of 250 million text-image pairs similar to JFT-300M [97] collected from the internet.
An extension of DALL-E known as DALL-E-2 [98] has also been presented, which can generate high-resolution images based on text prompts and existing images. It uses CLIP-based robust representation to enhance its image generation. DALL-E-2 first generates an image embedding induced by the text input using CLIP. Then, this generated image embedding is used as an input of the diffusion model to create an image. The model is trained using an image ($x$)-caption ($y$) pair. The proposed generative model comprises a prior ($P(z_{i}|y)$) producing CLIP image embeddings $z_{i}$ conditioned on $y$. Meanwhile, the decoder ($P(x|z_{i},y)$) generates images conditioned on CLIP image embeddings. These components are stacked to yield the generative model ($P(x|y)$).
4.4 Llama
LLaMa [99] is trained on publicly available datasets with over 1.4 trillion tokens and is based on existing large pre-trained models with key modifications. For instance, pre-normalization from GPT 3 was adopted to improve its training stability, but unlike in GPT-3, normalization was performed at the input of each transformer sub-layer. The SwiGLU [100] activation function from the PaLM model [82] replaced the ReLU non-linearity activation function. The absolute positional embeddings were replaced with rotary positional embeddings (RoPE) [101] from the PaLM model [82] at each layer. The training optimization included the use of the AdamW optimizer [102] and efficient implementation techniques for causal multi-head attention, backward passing, and memory usage reduction. LLaMA has 65B parameters several times more than GPT-3, and performed well in various downstream tasks. A more refined version called LLaMA 2 [103], with 70B parameters in addition to Llama 2-Chat, is designed for dialogue applications and performance surpasses open-source chat models on various benchmarks [103]. LLaMA2 incorporates increased context length and the implementation of grouped-query attention (GQA) to enhance inference scalability for larger models.
4.5 Stable Diffussion
Diffusion models (DMs) [104] can synthesize high-quality images but demand significant computational resources during training. Furthermore, they require costly evaluations in pixel space. Therefore, an efficient model called the stable diffusion model [105] explicitly separates the compressive and generative learning phases. The proposed model employs an autoencoder that generates a lower-dimensional representation space and significantly reduces computational complexity. This enhances the computational efficiency of sampling, leveraging the inductive bias of UNet-based diffusion models for spatially structured data and providing a multipurpose compression model applicable to various generative tasks. The proposed diffusion model demonstrates competitive performance in tasks such as unconditional image synthesis, inpainting, and super-resolution, significantly reducing computational costs compared to pixel-based DMs. An improved version of the stable diffusion model, known as SDXL [106], incorporates a larger UNet backbone (2.6B parameters in the UNet), additional attention blocks, and a broader cross-attention context. Furthermore, it uses a refinement model, enhancing visual fidelity through post-hoc image-to-image techniques.
4.6 Industrial pretrained models
Industrial-scale pre-trained models are extensively employed across a spectrum of industrial applications. Here we briefly mention some state-of-the-art models widely adopted in various industries including healthcare. LaMDA (Language Models for Dialog Applications) [107] is a transformer-based model with 137B parameters tailored for dialog applications. It addresses two challenges faced by language models: safety and factual grounding. The safety enhancements involve filtering responses using a classifier fine-tuned with crowd worker-annotated data, aligning model behavior with human values based on Google AI principles [^4] https://ai.google/responsibility/principles/. Factual grounding is achieved by enabling the model to consult external knowledge sources to ensure responses are grounded in verifiable information.
A more robust model than GPT-3 known as GLaM (Generalist Language Model) [108] has been proposed by Google, which has 1.2 trillion parameters. It utilizes sparsely activated mixture-of-experts (MoE) architecture [109] in which every other transformer layer replaces the feed-forward component with an MoE layer. GLaM achieved better performance on 29 NLP tasks in nearly all zero-shot, one-shot, and few-shot learning scenarios, consuming only one-third of the energy of GPT-3. A more refined version, the Pathways Language Model (PaLM) [82], with 540 billion parameters, has also been presented by Google. It is a decoder-only-style transformer model with multiple modifications, such as SwiGLU activations [100] for MLP intermediate layers, parallel formulation for a faster training speed, multi-query attention, and RoPE embeddings [101] for improved performance on long sequences. PaLM has achieved state-of-the-art performance in reasoning tasks and shows proficiency in multilingual tasks and source code generation. A more efficient model, PaLM 2 [110], was recently developed by Google and performs better than PaLM on various downstream tasks, with enhanced multilingual, reasoning, and coding capabilities. It is utilized at various sizes in 25 Google products, including Bard  [^5] https://bard.google.com/, the health-focused Med-PaLM [111], Med-Plam 2 [112], and the security-focused Sec-PaLM [^6] https://cloud.google.com/security/ai/. SenseNova [^7] https://www.sensetime.com/en/news-detail/51166397?categoryId=1072, developed by SenseTime, is a large-scale FM that offers various API interfaces and services specifically designed for enterprise customers. It is widely adopted in various industries, mainly in China, such as in the automotive sector, where SenseNova played a key role in the creation of the BEV (Bird’s-Eye-View) system that can recognize 3,000 types of objects. Furthermore, SenseNova introduced the “SenseChat” [^8] https://chat.sensetime.com/wb/login model, which is similar to ChatGPT but specifically designed for use in China.
Table 2: Flagship Foundation Models
Work Year Model Parameter Training Datasets Public Task
[16] 2018 BERTBASE 110M [93] [^9] BooksCorpus, English Wikipedia NLP
[16] 2018 BERTLARGE 340M [93] , English Wikipedia NLP
[17] 2019 RoBERTa 355M [93], English Wikipedia, CC-News, Opne Web Text [113], stories [114] NLP
[91] 2019 DistilBERT 66M [93], English Wikipedia NLP
[90] 2020 ALBERTBASE 12M [93], English Wikipedia NLP
[90] 2020 BERTLARGE 18M [93] , English Wikipedia NLP
[90] 2020 BERTxLARGE 60M [93], English Wikipedia NLP
[90] 2020 BERTxxLARGE 235M [93], English Wikipedia NLP
[15] 2018 GPT 117M [93] NLP
[60] 2019 GPT-2 1.5B WebText NLP
[29] 2020 GPT-3 175B Common Crawl (filtered), WebText2, Books1, Books2, Wikipedia NLP
[84] 2023 GPT-4 - NLP and image
[94] 2022 InstructGPT 1.3B Common Crawl (filtered), WebText2, Books1, Books2, Wikipedia NLP
[95] 2021 DALL-E 12B 250M image-text pairs from internet text-to-image synthesis
[98] 2022 DALL-E 3.5B 650M image-text pairs text-to-image synthesis
[99] 2023 LLaMA 65B CommonCrawl, C4 [19], Github, Wikipedia, Books, ArXiv, StackExchange NLP
[103] 2023 LLaMA2 70B 2 trillion tokens public data NLP
[105] 2022 Stable diffusion 1.45B LAION [115] Image synthesis
[106] 2023 SDXL 3.5B Not reported Image synthesis
[107] 2022 LaMDA 1.2T Dialog data, C4 Dialog Application
[108] 2022 GLaM 137B Filtered Webpages, Wikipedia, Conversations, Forums, Books, News NLP tasks
[82] 2022 PaLM 540b 780 billion tokens NLP
[110] 2023 PaLM2 13B NLP
[116] 2023 SenseNova - - NLP
Table 2: Flagship Foundation Models (continued)
5 Foundation Models in the Medicine Domain
Multiple FMs have been developed for general purposes that can be adopted across different domains. However, there has been a recent notable trend in the development of FMs, particularly within the healthcare domain [117]. These emerging models have shown to be highly successful (Table 3). This section highlights FMs exclusively designed for healthcare applications that tackle challenges and harness opportunities within the healthcare domain.
5.1 Clinical Large Language Models (CLLM)
BioBERT is one of the earliest domain-specific LLMs [118]. It is a biomedical language processing model that was trained on large-scale biomedical corpora. BioBERT demonstrated impressive performance across 15 datasets addressing three NLP tasks: named entity recognition (NER), relation extraction (RE), and question answering (QA). Another model called BioMegatron [119] employs a BERT-style architecture with up to 1.2 billion parameters. It is trained on a larger domain corpus (PubMed abstracts and PMC full-text) that outperforms previous state-of-the-art models on biomedical NLP benchmarks (NER, RE, QAs). GatorTron [120] is another state-of-the-art large-scale clinical language model developed from scratch and was designed for NLP tasks on EHRs. It has 8.9 billion parameters and is trained on a corpus of over 90 billion words. These data included a substantial amount of clinical text from the University of Florida (UF) Health, PubMed, Wikipedia, and MIMIC III [121]. It outperforms existing biomedical and clinical transformers in five clinical NLP tasks: clinical concept extraction, medical RE, semantic textual similarity, natural language inference, and medical QAs. ChatGPT is a general-purpose model trained on various datasets, including a portion of healthcare data. Although, the primary focus is not limited to the healthcare domain. Inspired by ChatGPT, GatorTron [120] was further extended to the GatorTronGPT [28] model. This was built on the GPT-3 architecture with 20 billion parameters. The training corpus included 277 billion words with 195 billion words from the Pile dataset and 82 billion words from de-identified clinical text obtained from UF Health. GatorTronGPT exhibits a remarkable ability to generate synthetic data (GatorTronS) aligned with medical knowledge, as validated through various clinical tasks. Furthermore, the performance on different tasks including NER, QA, RE, natural language inference, and semantic textual similarity achieved promising performance. Moreover, it has been tested on the physicians’ Turing test and achieved comparable performance to humans. Singhal et al. [111] proposed Med-PaLM, evaluated on a well-curated MultiMedQA benchmark. The proposed model achieved promising performance compared to existing models and was slightly inferior compared to clinicians. Another model built upon PaLM-2 (termed Med-PaLM-2) was fine-tuned on the MultiMedQA benchmark and achieved 19% better performance than Med-PaLM. The more recent MEDITRON model [122] is a large-scale medical language model with 70 billion parameters trained with 48.1 billion tokens from the medical domain. It has outperformed state-of-the-art models including GPT-3.5 and MedPaLM.
5.2 Medical Image Analysis
FMs are trained on vast numbers of training images that can be fine-tuned for various downstream tasks. However, the patterns and features of medical images differ significantly from those of natural images [123, 124]. Several attempts have been made to develop robust domain-specific models for medical images. For instance, the Segment Anything Model (SAM) [32] was originally developed for general-purpose segmentation and trained on over 11 million images. It has been fine-tuned for medical image applications with great success shown by [125]. A medical-related FM for universal medical image segmentation called MedSAM was proposed by Ma et al. [30] and works as a vision transformer-based (ViT) image encoder for feature extraction and a mask decoder for segmentation results. MedSAM uses multiple online medical datasets with 1,570,263 images and text pairs for training. A detailed evaluation showed promising performance compared to other segmentation models, including SAM [32] and U-Net [31]. Following the success of MedSAM, many models have been developed for medical image segmentation, including 3D medical image segmentation [126, 127, 128, 129]. The details of these segmentation tasks are explained in Section 6.2.3. Taher et al. [130] proposed Autodidactic Dense Anatomical Models (Adam), an SSL FM that learns human anatomy from medical images in a natural hierarchical manner. The proposed model consists of two modules: the anatomy decomposer (AD) and purposive pruner (PP). The AD decomposes anatomy into a hierarchy of anatomical structures, which guides the model to learn hierarchical relationships from the images. The PP ensures effective comprehension of anatomy by preventing semantic collision. This facilitates the learning of a wide range of distinct anatomical structures.
5.3 Text-to-Image
Text-to-image models aim to generate visually realistic images from textual descriptions and have emerged as one of the most impressive applications [131] in the era of FM. GAN-based works are able to generate images from text (see [131]). However,these models tend to suffer from lack of generalization, mode collapse, lack of diversity, and lack of delicacy in text-to-image details [10]. FMs such as DALL-E [95] can generate images for a wide range of textual descriptions. It also exhibits the ability to modify the attributes of objects and control the number of instances present in the generated images. Similarly, another model called Pathways Autoregressive Text-to-Image (Parti) [132] produces high-fidelity photorealistic images with complex compositions and rich content synthesis. However, these models suffer from high computation costs and sequential error accumulation [131]. Diffusion models [104, 105, 106], have emerged as the new state-of-the-art in text-to-image generation as a result. More information can be found in [131]. Text-to-image is also beneficial in healthcare such as work done by Kim et al. [133] who proposed FMs to generate text-conditional magnetic resonance imaging (MRI) scans. Similarly, Ali et al. [134] compared DALL-E 2, Midjourney, and Stable Diffusion to represent demographic disparities in the surgical profession. The authors analyzed images generated by each model using standardized prompts and geographic-based prompting. They compared them to real-world demographic data for attending physicians and trainees across eight surgical specialties. The findings reveal societal biases in Midjourney and Stable Diffusion, with a tendency to predominantly generate images of white male surgeons. DALL-E 2 presents a more balanced depiction of surgeons, although it still falls short of fully capturing diversity.
Figure 2: Learning architecture, model size, and training data used by representative foundation models.
5.4 Omics (Protein, Genomics, RNA)
Human cells are basic structural and functional units that contain different entities such as proteins, nucleic acids (DNA, RNA), carbohydrates, and lipids [135]. Multiple ML models have been used to extract useful patterns from the complex data structures associated with various omics domains, which include genomics, proteinomics, metabolomics, transcriptomics, Lipidomics, microbiomics, etc. [136]. However, due to lack of models generalizability, FMs are now widely used for omics data. For instance, Yang et al. [137] proposed a BERT-based approach named scBERT for Cell type annotation. The encoder of the BERT model cannot be scaled to more than 512 gene inputs, so it was replaced with a performer [138] that can be scaled to 1600 gene inputs. Thus, scBERT does not depend on dimensionality reduction techniques and is capable of the full gene-level interpretation. scBERT achieved better performance than existing models when evaluated on various tasks from seven scRNA-seq datasets. A BERT-based model for RNA named RNABERT [139] was proposed by Akiyama for ncRNA sequence analysis, focusing on structural alignment and clustering. Subsequently, other models for RNA-based data include 3’ untranslated regions [140] and 5’ untranslated regions [141] 101010 3’ and 5’ denote the ends of RNA or DNA molecules, with ”3’” representing the end where the third carbon atom of the sugar is located, and ”5’” representing the end where the fifth carbon atom is situated.. Chen et al. [142] showed that most of the existing models rely on annotated RNA sequences which results in small training data (less than 30K samples) [142] and omits large amounts of unannotated data. Therefore, they proposed a BERT-based RNA-FM model that utilizes 23 million unannotated RNA sequences in a self-supervised manner. However, Zhang et al. [143] showed that this approach struggles to capture evolutionary information from less conserved RNA sequences. Therefore, they proposed an RNA-MSM model [143] based on multiple sequence alignment using homologous sequences from the RNAcmap pipeline [144] Compared to RNA-MSM, RNA-FM performed worse than one-hot encoding in predicting base pairs and solvent-accessible surface areas. Another context-based deep learning model named Uni-RNA [145] was proposed for RNA sequences. Furthermore, a model named BigRNA [146] also utilizes RNA-seq data obtained from GTEx consortium V6 release [147] that can be applied to various downstream tasks for RNA biology. Chen et al. proposed SpliceBERT [148] for sequence-based RNA splicing prediction. It was trained on pre-mRNA (precursor messenger RNA) sequences from 72 vertebrates, creating a pre-training dataset of approximately 2 million sequences. A more recent model,scHyena, was proposed by Oh et al [149] for scRNA-seq analysis data in the brain. It incorporates a bidirectional Hyena operator [150], linear adaptor layer, and gene-based positional encoding to capture cellular diversity and outperforms existing methods in various tasks such as cell type classification and imputing biologically relevant values in scRNA-seq data.
Theodoris et al. [151] proposed the transformer-based Geneformer which is a context-aware model to map the gene networks. The Geneformer [151] architecture consists of six encoder units each with self-attention layers and a feedforward neural network, and is trained on 29.9 million single-cell transcriptomes (Genecorpus-30M collected from publicly available datasets). Cui [152] proposed the scGPT model trained on more than 33 million cells (from publicly available datasets) which captures meaningful insights from genes and cells. scGPT has been validated on a range of tasks such as multi-omics, call type annotations, clustering, batch correlation, genetic perturbation prediction, and gene network inference [152]. Yang et al. [153] developed GeneCompass, which is a cross-species transformer foundation model trained on scCompass-126M, which contains 120 million single-cell transcriptomes from humans and mice. Chen et al. [154] showed that although these models [151, 152, 137] achieved better performance, they heavily rely on well-curated datasets and are computationally expensive. Therefore, to address the challenges of content hallucination and answer genomics questions, GeneGPT was proposed, which utilizes ChatGPT embeddings of different genes from the literature. The authors created embeddings using GPT-3.5 from data of 33,000 genes obtained from Geneformer [151] and scGPT [152]. DNABERT [155] was one of the first few BERT-based models to decipher the language of non-coding DNA by capturing a global and transferrable understanding of genomic DNA sequences. An extension, DNABERT2 [156], replaced the k-mer tokenization with Byte Pair Encoding. It is trained on a multispecies dataset of 135 species from seven categories including fungi, protozoa, mammals, other vertebrates, and bacteria. It achieved promising performance with 21 times fewer parameters than DNABERT [155]. Cheng et al. [157] proposed the scGPT-based Gene Expression-based Molecule Generator (GexMolGen) for designing hit-like molecules from gene expression signatures. This model addresses the limitation of traditional similarity searching-based approaches, which search within a limited dataset. GexMolGen generates molecules capable of inducing the required transcriptome profile, exhibiting high similarity to known gene inhibitors. Fu et al. [158] proposed an interpretable foundation model named the General Expression Transformer (GET) for transcription regulation across 213 human fetal and adult cell types. GET achieves experimental-level accuracy in predicting gene expression, even in previously unobserved cell types. Fan et al. proposed a Genome Foundation Embedded Topic Model (GFETM) for the analysis of single-cell ATAC-seq (scATAC-seq) data. The GFETM consists of two components: the Embedded Topic Model (ETM) modeling the scATAC-seq data as a cells-by-peaks matrix, and Genome Foundation Models (GFMs) utilizing pre-trained models such as DNABERT and Nucleotide Transformers to extract peak nucleotide sequence embeddings. Another model, scBERT [137] addresses the challenges in annotating cell types using single-cell RNA-seq data. The scBERT model leverages pretraining on large unlabelled datasets followed by fine-tuning on specific data to achieve robust cell type annotation, novel cell type discovery, and interoperability. Alsabbagh et al. [159] evaluated the performance of scGPT [152], scBERT [137], and Geneformer [151] on an imbalance cell type dataset. The results on Zheng68K and Multiple Sclerosis show that scGPT [152] and scBERT [137] achieved better performance, but the Geneformer [151] performed the worst with 20% lower accuracy. The performance of scGPT and scBERT with logistic regression (LR) was evaluated in another study [160] for fine-tuning tasks in the cell type annotation. The authors show that LR achieved better or comparable performance to large-scale foundation models for RNA sequencing data analysis.
Table 3: Foundation Models in Healthcare domain
Work Year Model Parameter Dataset Dataset Size Public Task
[118] 2018 BioBERT 345M English Wikipedia, BooksCorpus, PubMed Abstracts, PMC Full-text articles 4.5B words Biomedical Text Mining
[119] 2020 BioMegatron 1.2B PubMed abstract set, CC0-licensed Commercial Use Collection of the PMC full-text corpus 6.1B words NER, QA, Relation extraction
[120] 2022 GatorTron 8.9B UF Clinical notes, PubMed articles, Wikipedia $>$90B words . NER, QA, Relation extraction, NLI, semantic textual similarity
[28] 2023 GatorTronGPT 20B UF Clinical notes, Pile dataset 277B words NER, QA, Relation extraction, NLI, semantic textual similarity
[111] 2023 Med-PaLM 540B MultiMedQA benchmark 780B tokens Medical QA
[112] 2023 Med-PaLM2 - MultiMedQA benchmark - Medical QA
[122] 2023 MEDITRON 70B Clinical Guidelines, PubMed, Experience Replay 48.1B tokens
[128] 2023 SAM-Med3D 100M 21K medical images 121K masks Medical 3d image segmentation
[127] 2023 SegVol 180M 25 CT datasets with 5772 CTs 90K unlabeled CTs image segmentation
[161] 2023 MultiTalent 69M 13 public abdominal CT datasets 1477 3D images image segmentation
[130] 2023 Adam 25M ChestX-ray14 [162] and EyePACS [163] $>$108K images Human organ identification.
[164] 2023 Endo-FM 121M 7 datasets for pre-training 5 million frames Endoscopy Video Analysis
[137] 2022 scBERT - Seven publicly available scRNA-seq datasets $>$ 1.8M cells Cell type annotation.
[153] 2023 GeneCompass 100M scCompass-126M (transcriptomes from human and mice) 120 M single-cell transcriptomes Cross-species genomics.
[152] 2023 scGPT 53M 8 datasets $>$ 10M cells Single-cell biology.
[139] 2022 RNABERT - SRNAcentral (human-derived small ncRNAs) 76237 ncRNAs ncRNA structural alignment and clustering.
[142] 2022 RNA-FM RNAcentral [165] $>$ 23M sequences RNA biology.
[143] 2023 RNA-MSM - rFam  [^11] https://rfam.org/ 4069 RNA families multiple sequence alignment.
[146] 2023 BigRNA GTEx consortium RNA biology.
[148] 2023 SpliceBERT 19.4M 2 million pre-mRNA sequences 2M RNA splicing prediction.
[154] 2023 GeneGPT 33000 genes genomics and single-cell transcriptomics.
[156] 2023 DNABERT2 117M 32.49 billion nucleotide from 135 species Genome FM.
[157] 2023 GexMolGen L1000 public dataset hit-like molecules, drug discovery.
[158] 2023 GET 80M single-cell Chromatin accessibility data 213 cell types Transcription regulation prediction.
[166] 2023 GFETM 2.5B GTEx consortium $>$ 10K cells scATAC-seq Analysis.
[149] 2023 scHyena GTEx consortium scRNA-seq analysis in the brain.
Table 3: Foundation Models in Healthcare domain (continued)
6 Medical Foundation Model Applications
6.1 Clinical NLP
6.1.1 Medical Chatbots and virtual assistants
FMs have advanced the use of chatbots in the medical domain. These medical chatbots are mainly based on BERT architecture and GPT-based models. For instance, Hamidi and Roberts [167] evaluated the performance of ChatGPT and Claude for patient-specific question answering (QA) from clinical notes and showed that they provided accurate and relevant answers. Furthermore, a study by Shahsavar and Choudhury [168] showed that the majority (78.4%) of participants (607 in total) were willing to use ChatGPT for self-diagnosis in healthcare. Similarly, medical chatbots can provide recommendations of comparable quality to those of human experts [168]. Huang et al. [169] showed that GPT-4 surpasses both GPT-3.5 and Family Medicine residents on multiple-choice questions from the University of Toronto Department of Family and Community Medicine Progress Test. Sun et al. [170] proposed PathAsst, a generative AI assistant for diagnostic and predictive analytics in pathology. PathAssist was trained on a 195K dataset of pathology image-text pairs and employed a base model architecture with a CLIP vision encoder and Vicuna-13B LLM. A voice-based digital assistance tool for ICU rounding teams was developed by King et al [171], that utilized real-time discussions to prompt evidence-based practices. FM-based chatbots in languages other than English have also been adopted [172, 173]. Incorporating FM-based medical chatbots and virtual assistants can improve healthcare accessibility and help relieve the burden on the healthcare system by facilitating communication between patients and healthcare providers. For instance, providing scalable and personalized patient support, augmenting clinical decision-making processes, and providing expert input to general practitioners when specialists are not available.
6.1.2 Content generation (clinical reports)
FMs have shown promising performance and assistance capability, especially in the generation of radiology reports. For instance, Clinical-BERT [174] can accurately predict disease from radiographs and generate reports. Similarly, RadioBERT [175], proposed by Kaur and Mittal, generates accurate medical reports from CXRs. Similar performance was achieved by Transformer-based Semantic Query (TranSQ). Beyond BERT-based approaches, including [176], GPT-based models have also achieved promising performance. For instance, Nakaura et al. [177] evaluated the performance of GPT models (GPT-2, GPT-3.5, GPT-4) in generating radiology reports based on patient information and imaging findings, and their performance was compared with reports generated by radiologists. Other studies [178, 179] have also shown that GPT-4 models can generate radiology reports with comparable quality to those of radiologists. Luo et al. [180] proposed a GPT-s-based model, BioGPT, that is specifically fine-tuned for various biomedical tasks, including end-to-end relation extraction and question answering. A study by Nicolson et al. [181] demonstrated the challenges associated with the diagnostic accuracy of CXR report generation. Their investigation focused on improving CXR report generation by warm-starting the encoder and decoder with recent ViT and LNP checkpoints. These advancements demonstrate the potential of FMs to improve diagnostic accuracy and assist radiologists in automating the reporting processes.
6.1.3 Information extraction
Xu et al. [182] proposed a BERT-based approach to link text phrases to ontology concepts, which can be employed in various downstream tasks including relation extraction and information retrieval. The proposed BERT-based approach achieved an accuracy of 83.5% in linking phrases to Unified Medical Language System (UMLS) concepts. Pan et al. [183] proposed a BERT-based medical text–to-SQL model (MedTS) to transform medical texts into SQL queries for Electronic Medical Records (EMRs) that achieved promising performance compared to existing techniques. Hanna et al. [184] evaluated the performance of ChatGPT in racial bias. The authors used ChatGPT to generate healthcare-related text for patients with HIV using 100 deidentified electronic health record encounters. The data was processed four times with variations in race/ethnicity while keeping other information constant. The generated text was then evaluated for sentiment, subjectivity, readability, NER, and word frequency. The results show that the differences in sentiment and subjectivity across races/ethnicities were not statistically significant. A BERT-based model was proposed by Kuling et al. [185] to segment breast radiology reports into different sections in the Breast Imaging Reporting and Data System (BI-RADS) lexicon which achieved 98% accuracy. FMs would enhance the efficiency and accuracy of extracting structured information from unstructured medical text data. Furthermore, such robust models offer promising avenues for improving data interoperability and accessibility in healthcare settings.
6.1.4 Information retrieval (bedside decision support)
Clinical decision support (CDS) plays a vital role in delivering information and tailored recommendations to patients and healthcare providers during care delivery. FMs can be adopted to significantly improve healthcare delivery. For example, Liu et al. [186] showed that ChatGPT could be a valuable complement in optimizing CDS alerts and advancing learning health systems. FM applications can facilitate accurate information retrieval to provide the most comprehensive and pertinent historical information to the provider.
6.1.5 Clinical Language translation
Medical language translation aims to transform complex medical texts such as those embedded in EHRs into a clearer format for diverse audiences [187]. This is achieved by simplifying complex medical terminology, abbreviations, and technical language so that they can be more easily understood by patients, caregivers, and individuals lacking specialized medical knowledge. For instance, “peripheral edema” or “hypertension” may be replaced with or tagged as “ankle swelling” or “high blood pressure”, respectively [187]. FMs show high accuracy for contextual medical language translation [187]. By accurately contextualizing medical terminology and simplifying complex language, FMs enable clearer and more accessible communication between healthcare providers, patients, and caregivers.
6.2 Medical Computer Vision
6.2.1 Generation (visual)
FMs have been widely adopted in the computer vision domain for multiple tasks, including content generation. For instance, Khader et al. [188] implemented diffusion probabilistic models to synthesize high-quality 3D MRI and CT images. The authors trained the VQ-GAN model on a dataset to obtain a low-dimensional latent representation suitable for training a 3D diffusion model. The results were validated by radiologists, proving that the model generates 3D volumetric medical imaging data that accurately reflects human anatomy. A stable diffusion-based model, RoentGen, for synthesizing chest X-rays (CXR) conditioned on text prompts was proposed by Delbrouck et al. [189]. The proposed RoentGen model can generate high-fidelity synthetic medical images with controllable features. Another work utilizing stable diffusion was carried out by Kidder [190], in which photorealistic medical images were generated including MRI, chest, and lung X-rays, and contrast-enhanced spectral mammography (CESM) images. Similarly, Moon et al. [191] proposed the Medical Vision Language Learner (MedViLL), which combines textual representations (BERT) and visual representations (with ResNet50 as the visual features extractor) for multiple downstream tasks such as diagnosis classification, medical image-report retrieval, and radiology report generation. Yang et al. [192] used the ImageCLEF 2023 dataset for medical report generation using BLIP-2 [193], with EVA-ViT-g as the encoder [194] and ChatGLM-6B as the decoder [195]. FMs could generate realistic data for training and validating algorithms thus reducing the use of expensive and sensitive real-world datasets. Furthermore, synthetic data can support the development of cutting-edge models without putting any patient information at risk [196].
6.2.2 Classification
DL models enable the prediction of predefined labels or categories associated with input data. The results are promising and often outperform medical experts in terms of accuracy [86]. Since DL-based models rely heavily on annotated datasets, they are not always suitable for generalization. Hence, FMs that utilize SSL or zero-shot classification can be a promising alternative. For instance, Tiu et al. [86] proposed CheXzero which demonstrated the effectiveness of SSL for pathology classification on CXR images without explicit annotations, achieving comparable accuracy to radiologists. Similarly, the efficient Global-Local Representations for Images using the Attention (Gloria) method [76] is an attention-based framework that contrasts image sub-regions and words in paired reports to global and local multimodal representations. Wang et al. [71] showed that decoupling image-text pairs effectively scale the training data and reduces the rate of false positives. However, models such as CheXzero and MedCLIP treat partial positive pairs as entirely negative pairs. Although MedCLIP reduces this issue, it still relies on expert labels for high performance. Therefore, Jang et al. [197] proposed a fine-tuning strategy by producing sentence sampling and positive-pair loss relaxation to achieve significant improvements (over 5%) in zero-shot pathology classification across various CXR datasets. You et al. [198] proposed CXR-CLIP, which utilizes image-text pairs with general prompts and multiple images and report sections. The proposed approach incorporated image and text contrastive losses to learn study-level characteristics of medical images and reports. A diffusion-based model, DiffMIC, was proposed by Yang et al. [199] to eliminate unexpected noise and perturbations in medical images and robustly capture semantic representations. The accuracy generalization of the classification models across various domains can be significantly improved by integrating techniques such as SSL, and attention-based frameworks. Furthermore, the development of diffusion-based models highlights the importance of robust semantic representation in medical imaging.
6.2.3 Segmentation
Medical image segmentation plays a vital role in medical diagnostics and provides sufficient non-invasive information to visualize, analyze, and track anatomical structures for accurate diagnosis and treatment planning [200]. Multiple works have carried out medical image segmentation tasks using deep learning models [200, 201, 202, 203, 204, 205, 206]. However, Recent ViT-based models include Swin UNEt TRansformers (Swin UNETR), proposed by Hatamizadeh et al. [207], which is a novel semantic segmentation model for brain tumor images that uses a sequence-to-sequence prediction approach. Swin UNETR combines a U-shaped network architecture with a Swin Transformer and has achieved state-of-the-art performance in the BraTS 2021 segmentation challenge. Other transformer-based models include TransUNET [208] and TransFuse [209]. Butoi et al. [210] showed that segmentation can be challenging due to the limited training data and the complex structures of human organs. Furthermore, DL-based models face challenges in generalization [211]. Therefore, they proposed a universal medical image segmentation (UniverSeg) model trained on a large dataset of over 22,000 scans collected from 53 open-access datasets from 16 different modalities. The proposed UniverSeg achieved promising performance compared to state-of-the-art models such as SENet [212], PANet [213], and ALPNet [214]. Kim et al. [215] showed that UniverSeg can be generalized for various tasks with minimal labeled data. Specifically, the authors demonstrated its promising performance for prostate MRI segmentation. Similarly, an application of SAM was evaluated for zero-shot tumor segmentation, tissue segmentation, and cell nuclei segmentation [125]. The performance of the SAM model was better for large connected objects but faced challenges in dense instance object segmentation. Another SAM-based model was also used for interactive medical image segmentation [216].
Xing et al. [217] showed that integrating denoising diffusion models (DDMs) into a U-shaped network architecture can improve the volumetric segmentation for multiple modalities. A DDM-based model was also proposed for dental radiography segmentation [218]. Another denoising diffusion probabilistic models (DDPM)-based [104] model that incorporated contextual discounting for tumor segmentation was used for multi-modality medical images. An improved diffusion model was proposed by Zhao et al. [219] that combines text attention principles with diffusion models for enhanced medical image segmentation precision. Wolleb et al. [220] introduced a novel semantic segmentation method for medical images that generates image-specific segmentation masks and computes pixel-wise uncertainty maps. The authors utilized diffusion models [221] using the BRATS2020 dataset to achieve meaningful segmentation masks that improved brain tumor segmentation performance compared to state-of-the-art models. Rahman et al. [222] proposed a diffusion-based segmentation approach to synthesize multiple variants of segmented masks based on various radiologist annotations. The authors proposed the Ambiguity Modelling Network (AMN) and Ambiguity Controlling Network (ACN) to capture uncertainty in ground truth and predicted masks. Furthermore, the Collective Insight (CI) score, combining sensitivity, the maximum Dice matching score, and the diversity agreement score was proposed. MedSegDiff, based on diffusion probabilistic models (DPMs), was proposed by Wu et al. [223] for segmentation in three modalities, such as optic cup segmentation from fundus images, brain tumor segmentation from MRI images, and thyroid nodule segmentation from ultrasound images. An extended version of MedSegDiff-V2 was proposed by Wu et al [224]. MedSegDiff-V2 integrated ViT into DPM and was evaluated for 20 segmentation tasks. Another diffusion model for vessel segmentation was proposed by Kim et al. [225] by combining the DDPM [104] with adversarial learning to diagnose vascular diseases. The proposed Diffusion Adversarial Representation Learning (DARL) model consists of a diffusion module to estimate latent features, a generation module for vessel segmentation masks and synthetic angiograms, and discriminators to distinguish between real and fake images. The results evaluated on multiple datasets achieved generalized and superior segmentation performance. Fu et al. [226, 227] presented denoising diffusion models that predict segmentation masks using a fully noise-corrupted input rather than relying on conventional ground truth masks during training. Sun et al. [228] proposed a diffusion-model-based method of automatic gland instance segmentation in histology images of the colon. Other interesting models and applications have been provided in [229, 228, 230, 231, 232, 233, 234, 235]. We outlined CLIP-based applications in Section 2.5, but some CLIP-driven applications specifically for image segmentation include those of [236]. These models showed that CLIP-based approaches exhibit reasonable transferability to medical images, especially in zero-shot settings for non-radiology photographic modalities. Similarly, Adhikari et al. [237] demonstrated the performance of CLIPseg and CRIS on synthetic data obtained from diffusion models. FMs can assist by enhancing diagnostic accuracy, surgical planning, disease staging, and medical decision-making for risk stratification. It can help to achieve faster diagnoses, tailored treatment decisions, and improved patient outcomes. Furthermore, incorporating FMs in clinical settings can assist clinicians in further refining segmentation accuracy and efficiency for improving the precision in various medical imaging applications.
6.2.4 Localization
FMs can serve as digital assistants to radiologists by drafting reports from imaging modalities which will substantially reduce their workloads [238]. Such automatically generated detailed reports can be used to identify abnormalities and contextualize findings based on the patient’s medical history [238]. Furthermore, FM-based grounded radiology reports can assist radiologists in providing more informed interactive conversations. Such as a radiologist might ask the model to identify new multiple sclerosis lesions compared to previous images [238]. Moreover, region-guided radiology report generation models enhance report completeness and consistency by focusing on individual anatomical regions and generating tailored descriptions [239]. Zhu et al. [240] proposed a novel visual grounding framework for multimodal fusion by aligning image-text embedding in a shared semantic space and utilizing a transformer-based deep interactor for accurate reasoning. However, implementation of such generalized models on a large scale may require the integration of data from multiple sources such as patient history, laboratory results, and previous imaging studies to provide contextually relevant insights [238]. Visual grounding from imaging modalities saves time and ensures consistency and completeness in reporting. Moreover, by contextualizing findings based on patient history, FMs facilitate more informed discussions between radiologists and patients.
6.2.5 Reconstruction
High-quality images are essential for clinical diagnosis, but addressing concerns related to cost and patient safety is also important [241]. Various imaging modalities, such as CT, MRI, and PET, require advanced reconstruction techniques to enhance spatial resolution, reduce noise, and improve overall image quality [242]. FMs have been shown to be effective for image reconstruction tasks. For example, Levac et al. [243] proposed deep generative diffusion models to solve the ill-posed inverse problem of accelerated image reconstruction with subject motion during MRI examinations. Similarly, Gungor proposed AdaDiff [244] for adaptive diffusion before MRI reconstruction. AdaDiff utilizes a diffusion process with an adversarial mapper for training and employs a two-phase reconstruction approach. The proposed approach is resilient against domain shifts in multi-contrast brain MRIs and has achieved promising performance for accelerated MRI reconstruction. The results from other works show that diffusion modeling is a promising approach for MRI reconstruction [245, 246]. FMs can improve healthcare accessibility by augmenting or enhancing image quality, particularly in facilities lacking advanced hardware or scanning capabilities and technical expertise.
6.2.6 Enhancement (super-resolution)
Image enhancement methods using super-resolution techniques can improve the quality and clarity of medical images. Super-resolution methods aim to reconstruct high-resolution images from lower-resolution inputs to enhance fine details. In medical imaging, they aim to provide precise visual information to clinicians [247]. Recently, multiple studies have been carried out on super-resolution techniques [247]. For instance, Mao et al. [248] proposed the Disentangled Conditional Diffusion Model (DisC-Diff) that incorporates a disentangled U-Net, forward and reverse diffusion processes, and novel loss functions to enhance the model’s effectiveness in capturing complex interactions in multi-contrast MRI. The proposed approach achieved promising performance compared to other works [249, 250]. Another robust latent diffusion-based model proposed in [251] showed promising performance in clinical MRI scans characterized by low resolution, diverse contrast, and spatial resolution variations. FMs significantly improve the quality and clarity of diagnostic images that enable clinicians to make more informed decisions that lead to better patient outcomes. For instance, FMs could allow hospitals with only 1.5T MRIs to synthesize 7T data, which would greatly affect clinical care and diagnostic accuracy.
6.2.7 Modality Translation
Multi-modality imaging such as CT-MRI provides key information for an improved decision-making process. However, obtaining multi-modality images can be expensive and requires acquisition time [252]. Therefore, synthesizing missing or corrupted modalities from available ones can provide a cost-effective medical image-to-image translation approach. Many DL-based models have been employed for this purpose, particularly GANs [253, 254, 255]. However, it has recently been shown that diffusion models are promising alternatives to GANs providing better results [256, 257]. For instance, SynDiff [256] employs a conditional diffusion process that progressively maps noise and source images onto the target image for efficient and high-fidelity translation between source and target modalities. Another diffusion-model-based approach utilizes 2D backbones for volumetric data synthesis [258]. FMs can be a promising approach to synthesizing multi-modality imaging and being able to translate information from one modality to another. This will assist clinicians with enhanced diagnostic capabilities and facilitate more informed treatment decisions.
6.2.8 Augmented procedures
Incorporating augmented reality (AR) into medical interventions and processes provides healthcare professionals with real-time, contextually relevant information to improve decision-making and overall patient care [259]. Augmented procedures can be applied across a spectrum of medical scenarios, including neurosurgery training, robotic neurosurgery, pain management, neuronavigation, rehabilitation, consent taking, and diagnostic tools [260]. By incorporating multi-modality data, FMs can potentially provide real-time assistance to surgical teams by assisting in various tasks including annotating video streams, alerting to skipped procedural steps, and diagnosing rare anatomical phenomena.
6.2.9 Vision-based conversation
Conversational agents (chatbots) gained significant attention, especially after ChatGPT in various domains including healthcare. Traditional chatbots in healthcare primarily relied on textual inputs from patients to aid and information. However, incorporating vision data can significantly enhance the capabilities of these conversational agents, enabling them to understand and respond to visual cues effectively. For instance, Luo et al. [261] proposed a multi-model FM, Valley (Video Assistant with Large Language model Enhanced ability), capable of understanding video, image, and language simultaneously. Qilin-Med-VL [262] is a Chinese large vision-language model for healthcare to comprehend both textual and visual medical data. Similarly, healthcare-based smart chatbots [263] intelligently respond to user input to detect conditions like viral and bacterial pneumonia in X-ray images and Drusen, Choroidal Neovascularization (CNV), and Diabetic Macular Edema (DME) in Optical Coherence Tomography (OCT) images. Furthermore, models such as VideoChat [264], VAST [265], visualGPT [266], NExT-GPT [267], CLIPSyntel [268] can aid in analyzing images or videos shared by patients to gather valuable insights into their physical condition, environment, symptoms to offer more personalized and context-aware support. FMs offer valuable assistance in analyzing visual information shared by patients to provide enhanced communication and diagnostic capabilities in healthcare settings.
6.3 Graph Learning
LLMs have excelled in natural language tasks, but they face challenges in graph learning due to their limitations in mathematical precision, logic reasoning, spatial perception, and the handling of temporal information. Therefore, work has been carried out to develop a framework specifically for graph data. Significant approaches include the incorporation of the capabilities of ChatGPT and the Toolformer [269] model for efficient graph reasoning [269]. Furthermore, incorporating multi-modality data provides better performance in graph learning models [270]. We refer the reader to [271, 272] for recent surveys on graph foundation models.
6.3.1 Brain network
Brain networks are complex and interconnected neural connections within the human brain [273]. Modern imaging technologies such as functional MRI (fMRI) have provided insight into normal and abnormal brain functions [273]. Graph neural networks (GNNs) show great potential in capturing complex interpretable brain network models for disease prediction [274]. Furthermore, GNN models are more common for investigating the brain connectome and how disruptions or alterations in brain networks are associated with conditions such as Alzheimer’s disease [275], Parkinson’s disease [276], psychiatric disorders [277], rare diseases [278], and autism spectrum disorders [279]. Due to their generalized nature, FMs offer a promising alternative to traditional GNNs for extracting meaningful patterns from brain data. FMs’ ability to capture complex interactions within brain networks makes them well-suited for deciphering neural connections associated with neurological disorders.
6.3.2 Knowledge discovery in medicine
Knowledge discovery in medicine also termed as hueristics is refers to the process of extracting valuable insights, patterns, and knowledge from large and complex medical datasets [280]. FMs are also adopted for knowledge discovery and drug discovery with great success [281]. Utilizing FMs in medical knowledge discovery can reveal novel correlations, biomarkers, and therapeutic targets for drug development, personalized medicine, and healthcare delivery.
6.4 Biology and Omics
FMs focused on biology and omics were outlined in Section 5.4. However, this section briefly explains the details of their applications. Additionally, we refer readers to several articles for detailed works concerned with the biology and omics domain: [282, 283, 284, 285, 286].
6.4.1 Drug discovery
Drug discovery aims to identify and develop new therapeutic compounds for treating various medical conditions. Traditional drug discovery methods have limitations due to small datasets and generalization challenges. However, the authors in[287, 288] demonstrated the potential of generative drug design and molecular foundation models. They achieved improved molecular quality and an enhanced capability for predicting drug properties. FMs could decrease the cost, time, manpower, and hardware requirements compared to traditional drug discovery and development. Furthermore, it also can impact high-throughput workflows.
6.4.2 Radiomics
Radiomics mainly involves the extraction and analysis of quantitative features from medical images to provide valuable insights into lesion characterization, prognosis, and treatment response [289]. Radiomics enables a more comprehensive and objective assessment of the underlying tissue properties by transforming conventional medical images (CT, MRI, PET) into high-dimensional data sets [290]. A work done by Pai et al. [291] proposed a SimCLR-based model [292] for managing biomarker discovery through self-supervised learning on a dataset comprising 11,467 radiographic lesions. The results show that FMs are efficient in learning imaging biomarkers and are accelerating their translation into clinical settings. Fu et al. [293] show that non-contrast CT-based radiomics can assist in identifying severe atherosclerotic renal artery stenosis (ARAS) patients suitable for percutaneous transluminal renal angioplasty (PTRA). FMs demonstrate remarkable potential in identifying imaging biomarkers and facilitating their translation into clinical practice. These capabilities have significant clinical implications such as better disease characterization, prognosis assessment, and personalized treatment strategies.
6.4.3 Proteomics
In 1995, Marc Wilkins created the term “proteome” for a new type of omics [294]. Proteomics is the study of the interactions, compositions, functions, and structures of proteins and their cellular activity [294, 295]. Many works have been carried out to develop intelligent solutions for proteomics[294, 295]. Some of the relevant FMs include [284, 285, 24, 283]. FMs can unlock new insights into complex cellular activities, paving the way for innovative approaches in drug discovery, personalized medicine, and disease management.
6.5 Other Applications
6.5.1 Text-to-Protein generation
Protein design aims to create customized proteins for various applications to address environmental and biomedical challenges. Leveraging recent advancements in transformer-based architectures has resulted in promising performance. For instance, leveraging contrastive learning and generative models, Liu et al. [296] proposed ProteinDT for text-guided protein generation to property prediction. Trained on a large-scale dataset (SwissProtCLAP, with 441K text and protein pairs), it outperformed state-of-the-art methods in protein property prediction benchmarks [297, 298, 299]. ProtGPT2 [284] was introduced as a language model that generates novel protein sequences following natural principles that exhibit amino acid propensities similar to natural proteins. ProGen [300] generates protein sequences with predictable functions across expansive protein families. Similarly, generalist medical AI (GMAI) [238] is a model capable of generating both amino acid sequences and three-dimensional structures of proteins based on textual prompts. Watson et al. [301] proposed a diffusion-based model for protein backbone design that fine-tunes the RoseTTAFold structure prediction network on protein structure denoising tasks. FMs present an innovative approach to generating protein sequences and structures from textual prompts. By integrating in-context learning capabilities, FMs could dynamically define tasks based on example instructions, such as generating proteins with specific binding properties.
6.5.2 Video
FMs were also successful in extracting insightful information from videos [302, 303, 261, 304]. For instance, Wang et al. [164] proposed the Endo-FM model for endoscopic video analysis based on a video transformer. The work used an SSL pre-training approach that captures spatial and temporal dependencies, pre-trained in a self-supervised manner on a large-scale dataset of 5 million frames obtained from 33K videos. A CLIP-based model, EchoCLIP [305] for echocardiography trained on a dataset of over a million cardiac ultrasound videos and expert interpretations achieved strong zero-shot performance in cardiac function assessment and identification of intracardiac devices. Other models proposed in[306, 307] are also large-scale pre-trained models for echocardiograms. FMs enable comprehensive understanding and analysis of dynamic visual data to enhance decision-making and diagnostic accuracy in healthcare.
6.5.3 Speech (interactive note taking)
Multiple works are done to utilize FMs for speech audio information. For instance, AudioGPT is based on the GPT model that integrates audio foundation models and input/output interfaces for spoken dialogue [308]. Similarly, various works show that FMs can complement the processing of speech data [309, 310, 311]. For instance, a study done by Gerczuk et al. [312] detects smaller temporal variations in depressed mood by introducing a zero-shot personalization strategy for large speech foundation models. FMs offer valuable capabilities for analyzing speech data and extracting meaningful insights that are not readily perceivable by healthcare providers.
Table 4: Applications of FMs in healthcare. CNLP: Clinical NLP, MCV: Medical computer vision
Work Domain Task Modality Dataset Pretrained Model Keypoints
[167] CNLP Evluation of ChatGPT and Claude EHR MIMIC III ChatGPT and Claude Accurate response
[168] CNLP ChatGPT for self-diagnosis Questionnaire Questionnaire ChatGPT Majority participants were satisfied from ChatGPT
[313] CNLP Chatbots evaluation against ophthalmologist Question-answer pairs Eye Care Forum (question-answer pairs) ChatGPT ChatGPT provide comparabable quality advice
[169] CNLP Chatbots (chatGPT 3.5, 4) evaluation against Family Medicine residents. Multiple-choice questions (MCQs) MCQs ChatGPT (3.5 and 4) GPT-4 performed the best.
[170] CNLP Diagnostic & Predictive Analytics Multimodal (image-text pairs) PathCap and PathInstruct [170] Vicuna-13B [314], CLIP (PLIP) [315] improved pathology diagnosis and treatment processes
[171] CNLP Digital assistant evaluation. Audio signals 106 patient audio recordings, PubMed Central manuscripts, and MIMIC III BioClinicalBERT FMs aid in reduced prompts per patients.
[182] CNLP concept normalization Text 1000 discharge summaries BERT Improved Accuracy
[183] CNLP Medical text to SQL queries EHR MIMICSQL BERT improved text-to-SQL generation
[188] MCV 3D MRI and CT Images MRNet Dataset, ADNI, Breast Cancer MRI Dataset, LIDC, IDRI, VQ-GAN DDPM High quality 3D volumetric data synthesis
[189] MCV chest X-ray (CXR) synthesis Multimodal (image-text pairs) MIMIC-CXR dataset Stable Diffusion High-fidelity synthetic CXRs conditioned on text.
[191] MCV Vision-Language Tasks Multimodal (image-text) MIMIC-CXR,VQA-RAD, and Open-I BERT, ResNet-50 Multi-modal understading of text-image pairs
[192] MCV Report generation Multimodal (image-text pairs) ImageCLEF 2023 BLIP-2, EVA-ViT-g as encoder, ChatGLM-6B as decoder FMs improved performance.
[190] MCV MRI, CXRs, CESM Multimodal (image-text pairs) [316, 317, 318, 319] Stable Diffusion generate photorealistic medical images
[184] CNLP Evaluate FMs for racial bias Text 100 EHR encounters ChatGPT No racial bias
[185] CNLP Section segmetation Text (radiology reports) 155,000 breast radiology reports BERT Imporved performance
[210] MMCV segmentation Images 53 datasets, 16 image modalities CNN Accurate segmentation
[215] MCV prostate MRI segmentation Images Prostate-Gland, PROSTATEx-TZ and PROSTATEx-PZ UniverSeg Can be generalized well.
[125] MCV Segmentation Images Tissue, TCGA, and nuclei [320] SAM [32] SAM face challenges to segment dense instance objects.
[321] MCV Segmentation, classification Images Ultrasound images (over 2 million) universal Ultrasound foundation model Modality specific.
[79] MCV Segmentation (25 organs, 6 tumors) Images 14 CT scans datasets CLIP Promising segmentation
[231] MCV Segmentation (organ/tumor segmentation) Images BTCV, LiTS [322] and JHH [323] CLIP Difficult to segment dense instance object
[220] MCV Generating segmentation masks Images Brats2020 DDPM Improved segmentation masks
[222] MCV Generating segmentation masks Images LIDC-IDRI, MSMRI, Bone surface segmentation AMN, ACN Improved segmentation masks
[225] MCV Generating segmentation masks Images XCAD, XCA, DRIVE, STARE DDPM, SPADE [324] Improved segmentation masks
[223] MCV Segmentation Images REFUGE-2 dataset [325], BraTs-2021 dataset [326] and DDTI dataset [327], DPM, ResUNet Improved segmentation masks
[224] MCV Segmentation Images AMOS2022, BTCV, REFUGE-2, BraTs-2021 dataset, ISIC 2018 dataset, and TNMIX dataset DPM, UNET segmentation
[328] MCV Segmentation Images LASeg dataset, Synapse dataset, MMWHS, M&Ms Diff-VNet [217] Improved segmentation
[227] MCV Segmentation Images abdominal CT[329], muscle ultrasound [330], Prostate MR [331], Brats2021 DDPM, DDIM Improved segmentation
[217] MCV Segmentation Images BraTS2020, MSD Liver, BTCV DDM, Diff-UNet Improved segmentation
[219] MCV Segmentation Images Kvasir-SEG [332], Kvasir-Sessile [333], and the GlaS [334] Diffusion model [104], DTAN Improved segmentation
[232] MCV Classification, Segmentation Image-text MS-CXR, MIMIC-CXR BERT, BioViL Improved performance
[174] CNLP Report generation Image-text MIMIC-III Clinical-BERT Diagnosis and report generation
[175] CNLP Report generation Image-text OpenI BERT, DistilBERT, RadioBERT Report generation from CXRs
[335] CNLP report generation Image-text IU X-ray, and MIMIC-CXR TranSQ medical report generation
[86] MCV X-ray classification Image-text MIMIC-CXR, CheXpert, PadChest CLIP SLL achieved better classification.
[71] MCV X-ray classification Image-text MIMIC-CXR, CheXpert,COVID, RSNA BioClinicalBERT, Swin Transformer SLL achieved better classification.
[197] MCV X-ray classification Image-text MIMIC-CXR, CheXpert, PadChest, Open-I, VinDr-CXR BERTbase, CLIP, SLIP, UniCLIP Improved pathology classification
[336] MCV classification Images PatchGastric CLIP, BioLinkBERT Better tubular adenocarcinoma classification
[198] MCV X-ray classification Image-text MIMIC-CXR, CheXpert, ChestX-ray14, RSNA pneumonia, SIIM Pneumothorax, VinDR-CXR, Open-I BioClinicalBERT, CLIP, ResNet50, Swin Tiny Improved pathology classification
[76] MCV image-text retrieval , Classification, Segmentation Image-text CheXpert, RSNA Pneumonia, SIIM Pneumothorax BioClinicalBERT, ResNet Improved performance
[199] MCV classification Images PMG2000, HAM10000, APTOS2019 DDPM, DCG model Improved performance
[245] MCV MRI reconstruction Images fastMRI [337], IXI  [^12] https://brain-development.org/ixi-dataset/ SSDiffRecon Improved reconstruction results.
[244] MCV MRI reconstruction Images fastMRI, IXI AdaDiff Improved reconstruction results.
[256] MCV MRI-CT translation Images IXI, BRATS, MRICT SynDiff Improved results
[258] MCV Modality translation Images S2M, RIRE , Make-A-Volume Improved results
[248] MCV MRI Super-Resolution Images IXI, in-house clinical brain MRI dataset DisC-Diff Improved results
[251] MCV MRI Super-Resolution Images IXI, UK Biobank. InverseSR Improved results
Table 4: Applications of FMs in healthcare (continued)
7 Opportunities
FMs present significant opportunities to advance healthcare by addressing key challenges and enhancing various aspects of medical practice. In this section, we explore three crucial challenges associated with FMs in healthcare.
7.1 Controllability
Controllability refers to the ability to guide and influence the outputs of FMs. FMs can be fine-tuned and customized to meet specific clinical requirements and references. Empowering clinicians with tools and interfaces to intervene in the model’s decision-making process will enable personalized and context-aware healthcare solutions. For instance, clinicians can use FM-generated recommendations based on patient-specific factors, clinical guidelines, and individual preferences to enhance the relevance and utility of FM outputs in clinical decision-making. Furthermore, controllability facilitates transparency for clinicians to understand the underlying reasoning of FMs. This will foster trust and confidence in their usage in medical practice.
7.2 Adaptability
Flexibility of FMs to adjust to diverse medical data, disease types, and healthcare contexts can be termed as adaptability. FMs are trained on large amounts of diverse medical datasets including medical literature, patient records, and clinical guidelines. FMs can be tailored and fine-tuned to accommodate various data modalities, disease phenotypes, and healthcare contexts. In practice, FMs can also provide real-time assistance through diagnosis, treatment planning, and decision-making. FMs also have the ability to continuously learn and adapt to new information by staying up-to-date with the latest information and practices. Harnessing such models will aid healthcare organizations in developing customized and context-aware solutions that align with specific clinical needs and operational requirements.
7.3 Applicability
Applicability refers to the flexibility of FMs in addressing a wide range of healthcare challenges and use cases. FMs can be applied to diverse tasks such as clinical decision support, medical image analysis, predictive modeling, natural language understanding, drug discovery, and precision medicine. Moreover, the applicability of FMs extends beyond traditional healthcare domains to encompass emerging areas such as digital health, telemedicine, population health management, and patient engagement. FMs are opening up new frontiers for innovation and transformation in healthcare delivery and patient care.
8 Challenges to FM in Healthcare
8.1 Cost
FMs are experiencing rapid growth with training parameters increasing significantly compared to traditional models. The models’ increasing size demands computationally expensive hardware requirements. This poses a handful of challenges in terms of the training cost, inference cost, debugging, monitoring, and maintenance in production applications [11].
8.2 Interpretability
FMs such as GPT-4 exhibit unparalleled capabilities for various tasks but execute them in unforeseen ways [11]. They have tremendous predictive capabilities beyond human imagination. However, it further complicates the efforts to understand their behavior. FMs undergo training across vast and diverse datasets of multiple domains and modalities, unlike task-specific models that are trained within specific domains. The extensive training results in the models learning a broad spectrum of behaviors. This makes it complex to distinguish the underlying mechanisms governing their performance across different tasks and domains [11]. Interpretability of FMs is challenging due to their vast training objective, complex structures, and task-agnostic capabilities. However, interpretability is still crucial as it ensures transparency, accountability, and trust for deploying such models in high-stake applications such as healthcare. Several works are now focusing on advancing the interpretability of FMs [338].
8.3 Validation
Validating the performance and efficiency of FMs in healthcare presents a significant challenge due to the complex and dynamic nature of medical data. Traditional validation approaches, such as cross-validation and holdout validation, may not be sufficient to assess the reliability and generalizability of FM outputs in real-world clinical settings. Moreover, the lack of standardized evaluation metrics and benchmarks tailored to healthcare applications complicates the validation process. One of the key challenges in validation is ensuring the robustness and stability of FM across diverse patient populations, disease cohorts, and clinical scenarios. Data quality, disease prevalence, and patient demographics can impact the generalization ability of FMs. This leads to biases, over-fitting, or under performance in specific contexts.
8.4 Verification
Verification of FMs in healthcare involves ensuring the correctness, reliability, and safety of model behavior and predictions. It involves various aspects such as model design, implementation, training, testing, and monitoring, to mitigate risks associated with erroneous or biased decision-making. One of the primary challenges in verification is validating the accuracy and consistency of FM outputs across different stages of the model life cycle. Ensuring reproducibility and consistency in model training, hyperparameter tuning, and optimization processes is essential to verify the reliability of FM-generated predictions. Additionally, verifying the robustness of FMs against adversarial attacks, data perturbations, and distribution shifts is critical to assess their resilience in real-world healthcare scenarios. Furthermore, verification should address concerns related to data privacy, security, informed consent, and algorithmic fairness to mitigate potential risks associated with FM usage.
8.5 Privacy and Security
Security and privacy considerations include protecting user data confidentiality, ensuring model integrity, and maintaining system availability. The possible threats to FMs include inference attacks, adversarial examples, and data poisoning [11]. Furthermore, since the FMs are adopted for various downstream tasks, they are prone to be compromised as a single point of failure. For instance, injecting malicious data as a training sample could impact all the dependent downstream applications [339].
8.6 Social bias
Social bias involves an uneven preference or bias towards one entity, individual, or group over another [340]. Unfairness arises when individuals are not treated equally in any tasks based on factors such as gender, race, age, and marital status [340, 341]. FMs may result in social bias [340, 342] such as significant social biases in Stable Diffusion [343], particularly in gender and racial stereotypes.
8.7 Scale
The volume of training data for FM development is expanding annually. This trend is expected to persist due to the reliance of models on publicly available datasets [344]. However, publicly available data is relatively smaller compared to industrial models [11]. Therefore, there is a need for models trained in large amount of data that are capable of handling the complexities associated with multi-modal foundation model datasets [345, 346].
8.8 Legal and ethical consideration
Legal and ethical considerations play a vital role in navigating the challenges posed by the deployment of FMs. From a legal perspective, issues arise on data collection methods, copyright laws, and privacy regulations [347]. Furthermore, regulations such as the General Data Protection Regulation (GDPR) influence the acquisition and utilization of training datasets. Another issue arises from the legal responsibility of model creators in cases when a model generates harmful responses, especially in healthcare [348]. Moreover, legal protections for model outputs may result in issues associated with free speech, ownership rights, and copyright laws [349]. In addition to the legal framework, ethical considerations are also important. Major ethical challenges associated with the development of FMs include obtaining informed consent for data usage, ensuring safety and transparency, addressing algorithmic fairness and biases, and safeguarding data privacy.
8.9 Accuracy/Performance
In healthcare, achieving high performance is important for the successful implementation of AI models. FMs present a promising foundation by being pre-trained on extensive and diverse datasets. However, they may not meet the specific accuracy requirements of healthcare especially due to hallucination  [^13] Hallucination refers to the generation of plausible but incorrect or unverified information. [350]. Fine-tuning these models with domain-specific training data is a common strategy to improve accuracy. However, this process can be resource-intensive. Striking a balance between accuracy improvement and resource allocation is crucial for effectively leveraging FMs in healthcare settings.
9 Discussion and Future Directions
FMs are powerful tools in various domains including healthcare that revolutionized medical practice and research. In this survey, FMs in medicine and their potential across multiple healthcare applications (Fig 1, Section 6) are explained. Most FMs are designed for general purposes to handle multiple tasks (Section 4). Such FMs exhibit the capacity to generalize across various tasks seamlessly, in a zero-shot manner. However, given the complex nature of healthcare data, applying such models directly may result in inaccurate inference. Therefore, these models are generally fine-tuned with more medical data [71]. Furthermore, FMs are also specifically designed for the healthcare domain (Section 5) including clinical LLMs (Section 5.1), medical image analysis (Section 5.2), and omics data (Section 5.4). Such models surpass the performance of general-purpose FMs when evaluated in medical data.
Most of the FMs are based on transformer architecture, which has been widely adopted after the great success since 2017 [44]. BERT was one of the first FMs designed for sequence modeling. FMs got attention after ChatGPT among the general public, leading to the development of similar models such as Bard, Chinese GPT, etc. FMs have revolutionized healthcare to assist clinical decision support, medical image analysis, drug discovery, and personalized medicine. Their ability to process large volumes of heterogeneous data, extract meaningful patterns, and generate insights has empowered healthcare professionals to improve diagnostic precision, personalized treatment strategies, and patient outcomes. For instance, FMs have been instrumental in developing medical chatbots, generating clinical reports, extracting information from medical texts, and facilitating language translation to enhance communication and information exchange in healthcare settings (Section 6.1). FMs also contributed to computer vision to facilitate various tasks including image generation, classification, segmentation, and modality translation (Section 6.2). FMs aid in reconstructing high-resolution medical images, enhancing image quality, and synthesizing missing or corrupted modalities. Moreover, such models have been integrated into augmented reality systems to enable real-time guidance and contextual information during medical interventions, training, and diagnostic procedures. Although FMs have tremendous potential in healthcare, however, several challenges need to be addressed to facilitate their widespread adoption and deployment (Section 7). Such challenges include cost implications, interpretability issues, validation and verification concerns, privacy and security risks, social biases, scalability limitations, and legal and ethical considerations. Although we explained the challenges in Section 7, addressing such challenges requires computational resources to train and deploy FMs. Moreover, the addressed model can be computationally expensive, especially for resource-constrained healthcare settings. Furthermore, ensuring the interpretability of FM outputs is crucial for fostering trust and transparency among clinicians and patients. However, interpreting the complex decision-making processes of FMs remains a challenge.
Fairness and Explainability of FMs: Healthcare is a sensitive domain in which false knowledge or hallucinations could lead to serious harm. For instance, the CLIP model’s sensitivity to red circles highlights how such biases may result in misclassifications [351]. Although researchers are trying to mitigate this issue, however, more research is required to improve the fairness of FMs and mitigate the potential risks of generating misinformation. Furthermore, incorporating the interpretability of FM decision-making processes will foster trust between the AI models and clinicians.
Privacy-preserving Techniques: Since most of the healthcare data is not publicly available, solving privacy issues of patient privacy and implementing federated learning [352] will provide FMs with a large amount of training data. Furthermore, it will enable collaborative model training across multiple institutions while ensuring data security and confidentiality.
Robust FMs: FMs that are susceptible to adversarial attacks may result in wrong information and diagnosis that leads to incorrect treatments or delays in necessary interventions [353]. Moreover, adversarial attacks can compromise patient privacy and may result in severe breaches of confidentiality. This would result in raising ethical and legal concerns. Although robustness has been slightly addressed recently [354, 355], there is still a need to incorporate robust FMs before deploying them into clinical settings.
Multi-purpose FMs: FMs are generally multi-purpose, however, due to the complex nature of healthcare data, general-purpose medical FMs are not yet explored that handle various applications. Designing a universal large-scale FM tailored specifically for healthcare would facilitate the creation of a centralized model that could be adopted across various healthcare domains.
Cost-effective Deployment: Developing efficient training and inference algorithms and exploring cloud-based solutions will facilitate cost-effective deployment in healthcare settings.
10 Conclusion
We conducted a detailed survey to review the recent advances of FMs in the healthcare domain. Our study explained the wide range of potential applications and significant contributions of FMs in revolutionizing various aspects of healthcare delivery. We provided insights into both the key general-purpose FMs and specialized models designed for healthcare applications. Moreover, our survey explored nearly all healthcare applications associated with FMs, ranging from clinical decision support systems to medical image analysis and drug discovery. By explaining the detailed landscape of the medical domain, our survey fills a crucial gap in existing literature and serves as a comprehensive resource for researchers and practitioners in the field.
11 Acknowledgement
This work was supported by the National Science Foundation (1908299, 2123809). A more compact version of this article is under view.